Simplify or solve if it is an equation.
213) SR2 + SR11 =

224) SR(6)*(SR6 – SR2) =

265) (SR5 + 3)(SR 2 – 8) =

Answer by Alan3354(69443)

Simplify or solve if it is an equation.
213) SR2 + SR11 =
224) SR(6)*(SR6 – SR2) =
265) (SR5 + 3)(SR 2 – 8) =
--------------
is SR a variable? Then you show SR(6) - is that a subscript?
Clarify things.